WHAT IS THE SLOPE FROM THE SET OF ORDERD PAIRS? (1,3) (-2,-4)
Found 3 solutions by mananth, stanbon, mahikab:
Answer by mananth(16946)

x1 y1 x2 y2
1 3 -2 -4

slope m = (y2-y1)/(x2-x1)
( -4 - 3 )/( -2 - 1 )
( -7 / -3 )
m= 7/ 3

Answer by stanbon(75887)

WHAT IS THE SLOPE FROM THE SET OF ORDERD PAIRS? (1,3) (-2,-4)
-------------
slope = (-4-3)/(-2-1) = -7/-3 = 7/3
=========================================
Cheers,
Stan H

Answer by mahikab(11)

The slope is 7/3.
The formula for slope is (y1-y2)/(x1-x2)
In this problem we can label the points like this:
x1 = 1
y1 = 3
x2 = -2
y2 = -4
Now we can input them into the formula:
(3-(-4))/(1-(-2))
= (3+4)/(1+2)
= 7/3
The slope is 7/3!